# Zero-Shot Tokenizer Transfer

### v1

@@ 1 Introduction
To close this gap, we introduce a new paradigm: We train a hypernetwork on a diverse distribution of tokenizers to predict the embedding parameters for any given tokenizer. By investing into the one-time-cost of training the hypernetwork, we aim to subsequently enable effective ZeTT. This proves to be possible: ZeTT via the hypernetwork preserves performance to a few percent accuracy in many cases. Furthermore, the hypernetwork can learn to rapidly adapt to a given target tokenizer by continued training on a small amount (<1B) of extra tokens, whereas previous work typically needed hundreds of billions of tokens (Dagan et al., 2024). As such, our hypernetwork provides a state-of-the-art solution to $n$-shot tokenizer transfer, while also establishing a competitive baseline to our newly introduced zero-shot tokenizer transfer problem. This unlocks a range of new ways to combine language models with tokenizers. For example, in this work, we zero-shot substitute the Mistral-7B tokenizer (Jiang et al., 2023) with a tokenizer that encodes code using 10% less tokens on average, while preserving functional code generation correctness to approx. 3% (Section 4.2). We also evaluate zero-shot cross-lingual transfer of the multilingual XLM-R encoder model to a range of different languages by substituting the XLM-R tokenizer with a target-language specific tokenizer and reusing adapters trained for the original XLM-R. This leads to a >16% speedup and preserves performance on XNLI (Conneau et al., 2018) to 1% on average, although the language model has never been trained with the target-language tokenizers. Finally, we show that a hypernetwork trained for a base large LM (e.g. Mistral-7B) can also be applied to fine-tunings of the same model (e.g. Mistral-7B-Instruct-v0.1), preserving capabilities to a large extent (Section 4.3). Our code and models are publicly available at github.com/bminixhofer/zett.
@@ 2 Background
Contemporary language models typically use subword tokenizers via BPE (Sennrich et al., 2016) or UnigramLM (Kudo, 2018). Subword tokenization is a common choice since it can represent arbitrary sequences of text ("open-vocabulary" language modeling) while largely retaining the efficiency of word-level models (Mielke et al., 2021). However, there are a number of problems with subword tokenization, e.g. models using subword tokenization struggle parsing sequences of numbers (Golkar et al., 2023) and text with spelling mistakes (Xue et al., 2022). A recent strand of work aims to get rid of subword tokenization via byte-level (so-called "token-free") models (Xue et al., 2022; Yu et al., 2023). However, these models still operate on tokens, using the set of 256 bytes as the vocabulary, and Unicode as the tokenization function (Mielke et al., 2021). In a similar vein, some models use character-level tokenization (Tay et al., 2022; Clark et al., 2022), optionally learning to pool characters into longer tokens (Nawrot et al., 2023). [^3] See also Mielke et al. (2021) for a comprehensive overview of tokenizers. So far, byte- or character-level approaches have been unable to supplant subword tokenization due to reduced compute efficiency (because of longer sequences), and not necessarily being more robust (Libovický et al., 2022). Thus, although our approach is in principle applicable to any tokenizer, we focus our experiments on subword tokenizers. Specifically, we use the UnigramLM parametrization of the tokenization function, and show that other tokenizers can be converted to this parametrization later in Section 5. UnigramLM sets
$T(x):=\operatorname*{argmax}_{C\in\mathcal{C}_{x}}\sum_{t\in C}\log p(t)$
where $\mathcal{C}_{x}$ is the set of all possible tokenizations of $x$ (i.e., all possible decompositions of $x$ in $\mathcal{V}$). UnigramLM provides a convenient way to represent tokens as a 2-tuple $(t,p(t))\in(\mathcal{V},\mathbb{R})$.
@@
Heuristic-Free Tokenizer Transfer. While a significant amount of prior work has investigated heuristics to initialize the embedding layer, there is also research into changing the training procedure to facilitate $n$-shot tokenizer transfer. Marchisio et al. (2023) show that forward- and backward-propagating through a subset of the model layers is sufficient for learning embeddings for a new tokenizer. Chen et al. (2023) find that regularly resetting the embedding parameters during pretraining boosts the speed at which they are relearnt upon transfer. These approaches can be seen as orthogonal to ours. They could be freely combined with our method; we leave this to future work.
Embedding Prediction Hypernetworks. Hypernetworks are networks that predict the parameters of another network (Ha et al., 2017). Prior work uses neural networks to predict embeddings for out-of-vocabulary (Pinter et al., 2017) or rare words (Schick & Schütze, 2019) of word embedding models (Mikolov et al., 2013). Schick & Schütze (2020) extend this approach to predict embeddings for rare words in BERT models (Devlin et al., 2019). These methods can also be viewed as embedding prediction hypernetworks. In contrast, the hypernetwork we propose (i) approaches the more general problem of transferring to an arbitrary tokenizer, instead of extending the original tokenizer and (ii) can be applied to encoder, decoder, and encoder-decoder LMs, that is, it is objective-agnostic.
@@ 3.1 Hypernetwork Training
3:     $\bm{q}\leftarrow\text{queue}(x_{1},..,x_{n}\sim\mathcal{D})$triangleright Create a pool of $n$ texts (where $n$geq $m$).
@@
7:         $\bm{q}\leftarrow\text{pop}(\bm{q},m)$triangleright Remove the least-recently-added batch.
8:         $\bm{q}\leftarrow\text{push}(\bm{q},x_{1},..,x_{m})$triangleright Add the current batch.
@@
10:         $\bm{t},\bm{f}\leftarrow\text{substrings}(\bm{q},l)$triangleright Compute all substrings and their frequency in $\bm{q}$.
11:         $\bm{f}\leftarrow\bm{f}/\sum_{i}f_{i}$triangleright Normalize frequencies to sum to one.
@@
14:              $p(t)\leftarrow f+\mathcal{N}(0,z^{2})$triangleright Assign a score based on frequency + noise to the substrings.
@@
16:         $\mathcal{V}_{b}\leftarrow\bm{t}[:k]$triangleright Assemble the top $k$ substrings into the tokenizer.
@@
19:         $\mathrm{loss}\leftarrow\mathcal{L}_{\theta}(T_{b}(\bm{x}),H_{\theta}(\mathcal{V}_{b},T_{b}),\psi)$triangleright Compute the loss on the $m$ texts in the current batch.
20:
21:         update $\theta$ using $\nabla\theta$ w.r.t. $\mathrm{loss}$.
@@
We initially fill a queue $\bm{q}$ with $n$ texts sampled randomly from the training corpus and, at every step in the training loop, push the $m$ texts in the current batch and remove the $m$ least recently added texts. We then compute all substrings $t$ up to length $l$ and their frequency in $\bm{q}$ [^4] In practice, implementing $\bm{q}$ as a queue allows efficiently caching the substrings and their probability $p(t)$ at this step. They only need to be recomputed for the new $m$ texts encountered in every batch. [^5] To ensure substrings do not cross word boundaries we pretokenize the text before computing substrings. We add Gaussian noise to the frequencies to arrive at a final score $p(t)$ for every token $t$. Finally, we assemble the tokenizer by taking the top $k$ tokens with the highest $p(t)$ as the vocabulary and UnigramLM parametrized by $p(t)$ as the tokenization function. The training loop is summarized in Algorithm 1. The ‘rolling’ queue of texts $\bm{q}$ ensures high variance in the vocabulary, while the Gaussian noise added to the frequencies ensures high variance in the tokenization function.
@@
$\mathcal{L}^{\text{warmup}}_{\theta}=\|H_{\theta}(\mathcal{V}_{a},T_{a})-\phi_{a})\|_{2}$
@@ 4 Experiments
Table 1: Accuracy on XNLI when reusing adapters trained for the original XLM-R model with new zero-shot transferred language-specific tokenizers. Also shown are the absolute change in accuracy from applying our hypernetwork ($\Delta$accuracy) and the average decrease in token length of the language-specific tokenizers over the original tokenizer ($\Delta$length).
@@ 4.1 Setup
Evaluation. We use the standard benchmarks PiQA (Bisk et al., 2020), HellaSwag (HS Zellers et al., 2019), BoolQ (Clark et al., 2019), MMLU (Hendrycks et al., 2021) and the “easy” subset of ARC (Clark et al., 2018) for evaluation in English and the synthesis task of HumanEvalPack (Muennighoff et al., 2023) for coding evaluation. For multilingual evaluation, we use XNLI (Conneau et al., 2018), XCOPA (Ponti et al., 2020) and MMLU as machine-translated by Lai et al. (2023).
Models. To evaluate our method, we use Mistral-7B (Jiang et al., 2023) as the main decoder-style language model and XLM-R (Conneau et al., 2020) as a representative of encoder-style models. [^9] Although (decoder-style) LLMs are the centerpiece of a large amount of current NLP research, encoder-style LMs have wide-ranging applications in e.g. retrieval (Khattab & Zaharia, 2020) and LLM distillation (Hsieh et al., 2023) due to their lower computational cost. We also experiment with the smaller TinyLlama-1.1B model (Zhang et al., 2024) in Appendix G.
@@
original FOCUS ours $\Delta$accuracy $\Delta$length
@@
Hypernetwork training. We train the hypernetwork for 200k gradient update steps (10k of which are MIMICK-style warmup) with a batch size of 128 tokens and a sequence length of 128 (we find it sufficient to use short sequence lengths for learning embedding parameters). For the multilingual decoder-style models, we start from the English + Code checkpoint and forgo MIMICK-style warmup, keeping other hyperparameters unchanged. We use a RoBERTa-style architecture i.e. bidirectional attention and Post-LayerNorm Transformer layers (Liu et al., 2019), but use a feedforward dimension of 2x the hidden dimension (instead of RoBERTa’s 4x) for the hypernetwork. See Appendix D for a full list of hyperparameters.
@@ 4.2 Zero-Shot and n-shot Results
Results for XLM-R are shown in Table 1. We take task adapters trained for the original XLM-R model on the English XNLI dataset via Poth et al. (2023) and substitute the tokenizer for our language-specific one. We compare our hypernetwork against a simple lexical baseline (copying the embeddings of overlapping tokens and initializing the rest randomly), FVT, OFA, and FOCUS (c.f. Section 2). We focus only on FOCUS in the following since it performs best among the baselines. [^11] We note, however, that FVT comes close to FOCUS’ performance without requiring auxiliary embeddings so it may be a better choice for practical applications. Our hypernetwork consistently outperforms all baselines and preserves accuracy to 1% on average, losing 3% in the worst case and improving by 1% in the best case, while sequences are on average 14% shorter for the language-specific tokenizers; inference is thus more than 16% faster.1212121/(1-14%)=16%, plus additional speedup due to attention scaling quadratically with sequence length. We show in Appendix E that these results are robust to the target vocabulary size.
@@ 4.3 Applying a Hypernetwork trained for a Base Model to Fine-Tuned Models
So far, we have shown that the hypernetwork can be successfully applied for transferring the tokenizer of the base model [^13] We refer to models purely pretrained on the Language Modeling task as base models. it was trained on. However, a large amount of the models used by practitioners are fine-tuned versions of base models, e.g. via SFT or RLHF (Ouyang et al., 2022). We now attempt to answer the question: Given a hypernetwork trained for a base model, can we apply this hypernetwork to fine-tuned versions of the same model without any extra training? This would act as a multiplying factor for the hypernetwork’s applicability. First, we observe that the embedding space of a fine-tuned model is compatible with that of the base model: the embeddings of the fine-tuned Mistral-7B-Instruct-v0.1 have an average cosine similarity of 98.6% to the corresponding embedding in the base model while the average cosine similarity of the mean embedding vector is 17.4%. [^14] Averaged across the input and the output embeddings. Embedding compatibility also holds true for other models (Appendix G). The predictions of a hypernetwork trained for a base model can thus be used out-of-the-box with fine-tuned models. We verify that this is the case by evaluating Mistral-7B-Instruct-v0.1 transferred to the GPT2 tokenizer on the corrected [^15] Using the corrections from https://github.com/InflectionAI/Inflection-Benchmarks. version of MT-Bench (Zheng et al., 2023). For $n$-shot transfer, since we train the full model we also need a way to transfer the non-embedding parameters; we achieve this via Task Arithmetic (Ilharco et al., 2023). Results are shown in Table 5.
The transferred fine-tuned model performs well, coming within approx. 0.5 score of the original model. Also, curiously, the fine-tuned model with the original tokenizer performs better when using the embeddings of the (not fine-tuned) base model; this may be a prudent direction for future work.
@@ 5 Discussion
BERT Mistral-7B TinyLlama-1.1B GPT2
Kind WordPiece BPE BPE BBPE
Original $p(\text{preserved})$ 100% 100% 100% 100%
bits per char n/a 0.675 0.747 0.930
To Byte-Level $p(\text{preserved})$ 99.6% 99.9% 99.9% 100%
Extra Tokens 162 522 362 0
Unigramify $p(\text{preserved})$ 99.4% 99.8% 99.8% 99.7%
bits per char n/a 0.678 0.750 0.932
Model Embeddings Tokenizer $\bm{(\mathcal{V},T)}$
@@
GPT2 original 0.930 0.932 1.005
ours 0.919 0.920 0.964
ours (no noise) 0.925 0.926 0.978
Converting tokenizers to byte-level. As per Section 3.2, we need a procedure to convert tokenizers to the byte level to ensure that token decomposition is always possible. This is trivial in most cases; the bytes just need to be added to the vocabulary. BPE is an exception: here, we need to change the atomic units on which merges are defined from characters to bytes. This can be achieved by adding merges to assemble the characters used by the tokenizer from their constituent bytes to the beginning of the merge table. We measure the success of the conversion to byte level as the probability that, given some pretoken sampled from a corpus, this pretoken results in the same token sequence in the original and the converted tokenizer. Results are shown in Table 6.
Converting tokenizers to UnigramLM. We also introduce a procedure to convert arbitrary tokenizers to tokenizers using UnigramLM as the tokenization function. We refer to this process as unigramifying (details in Appendix A). An important assumption of the hypernetwork training is that by using the UnigramLM parametrization with scores distributed as Gaussians we can cover a sufficiently diverse distribution of tokenizers to enable the hypernetwork to generalize to e.g. BPE tokenizers. Unigramifying allows us to check if, in principle, this is possible. Luckily, we find that it is: unigramifying results in minimal performance degradation when substituting the original tokenizer with the corresponding UnigramLM tokenizer (Table 6). Although this does not guarantee that our distribution of tokenizers is sufficiently diverse, our empirical results suggest it is (cf. Section 4.2).
@@
What is the effect of amortizing over the tokenization function? As described earlier in Section 3, we ‘amortize’ over the tokenization function, that is, the tokenization function is not an input to our hypernetwork. We find that the predicted amortized embeddings are robust to the choice of tokenization function. For example, the set of embeddings predicted for the GPT2 vocabulary has low bits-per-character for both the original GPT2 tokenization function and a different UnigramLM tokenization function with scores based on token frequencies (Table 7). This is not the case for the original GPT2 embeddings: while they (as expected) perform well with the original GPT2 tokenizer, there is significant performance degradation when switching to the frequency-based UnigramLM tokenization function. This calls into question prior work copying the embeddings of overlapping tokens for transfer across tokenizers (Dobler & de Melo, 2023; Gee et al., 2022, among others), indicating that even if there is an exactly overlapping token in the original tokenizer, it is not necessarily the optimal initialization of the corresponding token in the new tokenizer.
@@
Although we amortize over most of the aspects of the tokenization function, in practice, tokenization functions rely on a considerable amount of engineering, so it is not possible to amortize over everything; we discuss remaining assumptions in Appendix H.
Model Hypernet
#params FLOPs / token #params FLOPs / token
GPT2 124M 253M 21M (16%) 4.5M (1.8%)
TinyLlama-1.1B 1.1B 2.1G 170M (15%) 33.1M (1.6%)
Mistral-7B 7.2G 15.4G 678M   (9%) 132.1M (0.9%)
@@
Analyzing computational overhead of the hypernetwork. We estimate the FLOPs per token of multiple hypernetworks in Table 8.161616We estimate FLOPs on the basis of XLA-compiled instructions using Jax (Bradbury et al., 2018). Given a batch size $n$ and sequence length $s$ for the main model, and using the hypernetwork to compose $k$ token sequences of length $t$, the FLOPs per batch will be $n\times s\times(\frac{\text{FLOPs}}{\text{token}})_{\text{main}}+k\times t\times(\frac{\text{FLOPs}}{\text{token}})_{\text{hypernet}}$. Taking hypernet training for Mistral-7B as an example with $n=s=128$, $k=32768$ and $t=7$ the FLOPs per batch will be 252T + 30T i.e. a 12% overhead from applying the hypernet. Notably, we observed in preliminary experiments that a hypernetwork size of three layers is sufficient, regardless of model size, so the relative overhead decreases with increased amounts of layers in the main model (as also evident in Table 8).

### v2

@@ 1 Introduction
To close this gap, we introduce a new paradigm: We train a hypernetwork on a diverse distribution of tokenizers to predict the embedding parameters for any given tokenizer. By investing in the one-time cost of training the hypernetwork, we aim to subsequently enable effective ZeTT. This proves to be possible: ZeTT via the hypernetwork preserves performance to a few percent accuracy in many cases. Furthermore, the hypernetwork can learn to rapidly adapt to a given target tokenizer by continued training on a small amount (<1B) of extra tokens, whereas previous work typically needed hundreds of billions of tokens (Dagan et al., 2024). As such, our hypernetwork provides a state-of-the-art solution to $n$-shot tokenizer transfer, while also establishing a competitive baseline for our newly introduced zero-shot tokenizer transfer problem. This unlocks a range of new ways to combine language models with tokenizers. For example, in this work, we zero-shot substitute the Mistral-7B tokenizer (Jiang et al., 2023) with a tokenizer that encodes code using 10% fewer tokens on average, while preserving functional code generation correctness to approx. 3% (Section 4.2). We also evaluate zero-shot cross-lingual transfer of the multilingual XLM-R encoder model to a range of different languages by substituting the XLM-R tokenizer with a target-language specific tokenizer and reusing adapters trained for the original XLM-R. This leads to a >16% speedup and preserves performance on XNLI (Conneau et al., 2018) to 1% on average. Finally, we show that a hypernetwork trained for a base large LM (e.g. Mistral-7B) can also be applied to fine-tunings of the same model (e.g. Mistral-7B-Instruct-v0.1), preserving capabilities to a large extent (Section 4.3).
@@ 2 Background
Contemporary language models typically use subword tokenizers via BPE (Sennrich et al., 2016) or UnigramLM (Kudo, 2018). Subword tokenization is a common choice since it can represent arbitrary sequences of text ("open-vocabulary" language modeling) while largely retaining the efficiency of word-level models (Mielke et al., 2021). However, there are a number of problems with the (lack of) robustness of subword tokenization (Xue et al., 2022; Golkar et al., 2023). A recent strand of work aims to get rid of subword tokenization via byte-level (so-called "token-free") models (Xue et al., 2022; Yu et al., 2023). However, these models still operate on tokens, using the set of 256 bytes as the vocabulary, and UTF-8 as the tokenization function (Mielke et al., 2021). In a similar vein, some models use character-level tokenization (Tay et al., 2022; Clark et al., 2022), optionally learning to pool characters into longer tokens (Nawrot et al., 2023). So far, byte- or character-level approaches have been unable to supplant subword tokenization due to longer sequences resulting in higher compute requirements, and not necessarily being more robust (Libovický et al., 2022). Thus, although our approach is applicable to any tokenizer, we focus our experiments on subword tokenizers. Specifically, we use the UnigramLM parametrization of the tokenization function, and show that other tokenizers can be converted to this parametrization later in Section 5. UnigramLM sets $T(x):=\operatorname*{argmax}_{C\in\mathcal{C}_{x}}\sum_{t\in C}\log p(t)$ where $\mathcal{C}_{x}$ is the set of all possible decompositions of $x$ in $\mathcal{V}$. This provides a convenient way to represent tokens as a 2-tuple $(t,p(t))\in(\mathcal{V},\mathbb{R})$.
@@
Heuristic-Free Tokenizer Transfer. In addition to heuristics, there is also research into changing the training procedure to facilitate $n$-shot tokenizer transfer. Marchisio et al. (2023) show that forward- and backward-propagating through a subset of the model layers is sufficient for learning embeddings for a new tokenizer. Chen et al. (2023) find that regularly resetting the embedding parameters during pretraining boosts the speed at which they are relearnt upon transfer. These approaches can be seen as orthogonal to ours. They could be freely combined with our method; we leave this to future work.
Embedding Prediction Hypernetworks. Hypernetworks are networks that predict the parameters of another network (Ha et al., 2017). Prior work uses hypernetworks to predict embeddings for out-of-vocabulary (Pinter et al., 2017) or rare words (Schick & Schütze, 2019, 2020) of word embedding models (Mikolov et al., 2013) and BERT (Devlin et al., 2019). In contrast, our hypernetwork (i) approaches the more general problem of transferring to an arbitrary tokenizer, instead of extending the original tokenizer and (ii) can be applied to encoder and decoder LMs, that is, it is objective-agnostic.
@@ 3.1 Hypernetwork Training
3:  $\bm{q}\leftarrow\text{queue}(x_{1},..,x_{n}\sim\mathcal{D})$ $\triangleright$ Create a pool of $n$ texts (where $n$ $\geq$ $m$).
@@
7:    $\bm{q}\leftarrow\text{pop}(\bm{q},m)$ $\triangleright$ Remove the least-recently-added batch.
8:    $\bm{q}\leftarrow\text{push}(\bm{q},x_{1},..,x_{m})$ $\triangleright$ Add the current batch.
@@
10:    $\bm{t},\bm{f}\leftarrow\text{substrings}(\bm{q},l)$ $\triangleright$ Compute all substrings and their frequency in $\bm{q}$.
11:    $\bm{f}\leftarrow\bm{f}/\sum_{i}f_{i}$ $\triangleright$ Normalize frequencies to sum to one.
@@
14:     $p(t)\leftarrow f+\mathcal{N}(0,z^{2})$ $\triangleright$ Assign a score based on frequency + noise to the substrings.
@@
16:    $\mathcal{V}_{b}\leftarrow\bm{t}[:k]$ $\triangleright$ Assemble the top $k$ substrings into the tokenizer.
@@
19:    $\mathrm{loss}\leftarrow\mathcal{L}_{\theta}(T_{b}(\bm{x}),H_{\theta}(\mathcal{V}_{b},T_{b}),\psi)$ $\triangleright$ Compute the loss on the $m$ texts in the current batch.
20:    update $\theta$ using $\nabla\theta$ w.r.t. $\mathrm{loss}$.
@@
We initially fill a queue $\bm{q}$ with $n$ texts sampled randomly from the training corpus and, at every step in the training loop, push the $m$ texts in the current batch and remove the $m$ least recently added texts. We then compute all substrings $t$ up to length $l$ and their frequency in $\bm{q}$. [^3] In practice, implementing $\bm{q}$ as a queue allows efficiently caching the substrings and their probability $p(t)$ at this step. They only need to be recomputed for the new $m$ texts encountered in every batch. [^4] To ensure substrings do not cross word boundaries we pretokenize the text before computing substrings. We add Gaussian noise to the frequencies to arrive at a final score $p(t)$ for every token $t$. Finally, we assemble the tokenizer by taking the top $k$ tokens with the highest $p(t)$ as the vocabulary and UnigramLM parametrized by $p(t)$ as the tokenization function. The training loop is summarized in Algorithm 1. The ‘rolling’ queue of texts $\bm{q}$ ensures high variance in the vocabulary, while the Gaussian noise added to the frequencies ensures high variance in the tokenization function.
@@
$\mathcal{L}^{\text{warmup}}_{\theta}=\|H_{\theta}(\mathcal{V}_{a},T_{a})-\phi_{a}\|_{2}$
@@ 4 Experiments
[…]
@@ 4.1 Setup
Evaluation. We use the standard benchmarks PiQA (Bisk et al., 2020), HellaSwag (HS; Zellers et al., 2019), BoolQ (Clark et al., 2019), MMLU (Hendrycks et al., 2021) and the “easy” subset of ARC (Clark et al., 2018) for evaluation in English and the synthesis task of HumanEvalPack (Muennighoff et al., 2023) for coding evaluation. For multilingual evaluation, we use XNLI (Conneau et al., 2018), XCOPA (Ponti et al., 2020) and MMLU as machine-translated by Lai et al. (2023).
Models. To evaluate our method, we use Mistral-7B (Jiang et al., 2023) as the main decoder-style language model and XLM-R (Conneau et al., 2020) as a representative of encoder-style models. [^8] Although (decoder-style) LLMs are the centerpiece of a large amount of current NLP research, encoder-style LMs have wide-ranging applications in e.g. retrieval (Khattab & Zaharia, 2020) and LLM distillation (Hsieh et al., 2023) due to their lower computational cost. We also experiment with the smaller TinyLlama-1.1B model (Zhang et al., 2024) in Appendix H.
@@
original FOCUS ours $\Deltaaccuracy Δ\Deltalength$
@@
Hypernetwork training. We train the hypernetwork for 200k steps (10k of which are MIMICK-style warmup) with a batch size of 128 and a sequence length of 128 (we find it sufficient to use short sequence lengths). [^10] Training takes around one day for the XLM-R hypernetwork on a TPU v3-8 and three days for the Mistral-7B hypernetwork on a TPU v4-32 pod. For the multilingual decoder-style models, we start from the English + Code checkpoint and forgo MIMICK-style warmup, keeping other hyperparameters unchanged. We use a RoBERTa-style architecture i.e. bidirectional attention and Post-LayerNorm Transformer layers (Liu et al., 2019), but use a feedforward dimension of 2x the hidden dimension instead of 4x for the hypernetwork. See Appendix D for a full list of hyperparameters.
@@ 4.2 Zero-Shot and n-shot Results
Results for XLM-R are shown in Table 1. We take task adapters trained for the original XLM-R model on the English XNLI dataset via Poth et al. (2023) and substitute the tokenizer for our language-specific one. We compare our hypernetwork against a simple lexical baseline (copying the embeddings of overlapping tokens and initializing the rest randomly), FVT, OFA, and FOCUS (c.f. Section 2). We focus only on FOCUS in the following since it performs best among the baselines. Our hypernetwork consistently outperforms all baselines and preserves accuracy to 1% on average, losing 3% in the worst case and improving by 1% in the best case, while sequences are on average 14% shorter for the language-specific tokenizers; inference is thus more than 16% faster.1111111/(1-14%)=16%, plus additional speedup due to attention scaling quadratically with sequence length. We show in Appendix E that these results are robust to the target vocabulary size.
@@ 4.3 Applying a Hypernetwork trained for a Base Model to Fine-Tuned Models
A large amount of the models used by practitioners are fine-tuned versions of base models [^12] We refer to models purely pretrained on the Language Modeling task as base models., e.g. via SFT or RLHF (Ouyang et al., 2022). We now attempt to answer the question: Given a hypernetwork trained for a base model, can we apply this hypernetwork to fine-tuned versions of the same model without any extra training? This would act as a multiplying factor for the hypernetwork’s applicability. First, we observe that the embedding space of a fine-tuned model is compatible with that of the base model: the embeddings of the fine-tuned Mistral-7B-Instruct-v0.1 have an average cosine similarity of 98.6% to the corresponding embedding in the base model while the average cosine similarity of the mean embedding vector is 17.4%. [^13] Averaged across the input and the output embeddings. Embedding compatibility also holds true for other models (Appendix H). The predictions of a hypernetwork trained for a base model can thus be used out-of-the-box with fine-tuned models. We verify that this is the case by evaluating Mistral-7B-Instruct-v0.1 transferred to the GPT2 tokenizer on the corrected [^14] Using the corrections from https://github.com/InflectionAI/Inflection-Benchmarks. version of MT-Bench (Zheng et al., 2023). For $n$-shot transfer, since we train the full model we also need a way to transfer the non-embedding parameters; we achieve this via Task Arithmetic (Ilharco et al., 2023). Results are shown in Table 5. The transferred fine-tuned model performs well, coming within approx. 0.5 score of the original model. Also, curiously, the fine-tuned model with the original tokenizer performs better when using the embeddings of the (not fine-tuned) base model; this may be a prudent direction for future work.
@@ 5 Discussion
Table 6: NLI performance on Farsi (FarsTail; Amirkhani et al., 2023), Dutch (SICK-NL; Wijnholds & Moortgat, 2021), Aymara and Guarani (AmericasNLI; Ebrahimi et al., 2022). We measure zero-shot transfer from a model trained on English XNLI (c.f. Table 1), except for Sick-NL where we train an adapter on SICK (Marelli et al., 2014) since the XNLI adapter underperforms.
Unseen by Hypernet Completely Unseen
Farsi Dutch Aymara Guarani
original 72.4 76.6 40.0 42.4
Lexical 60.5 72.7 38.5 41.7
FVT 65.5 74.8 38.4 39.0
FOCUS 65.3 74.8 37.7 40.7
ours 66.4 77.8 42.9 42.2
@@
[…]
Table 7: Performance of Mistral-7B transferred to the GPT2 tokenizer on English benchmarks (c.f. Table 2), as well as transferred to a tokenizer containing all words in the evaluation datasets; this converts Mistral-7B to a word-level language model on the evaluation corpora.
PiQA HS ARC BoolQ MMLU Avg.
original 80.7 81.0 79.5 83.6 59.6 76.9
GPT2 Tokenizer FOCUS 69.2 63.8 45.7 60.4 38.8 55.6
ours 79.7 77.5 73.0 81.9 53.0 73.0
[…]
Word Tokenizer FOCUS 66.8 58.8 51.3 62.6 35.2 54.9
ours 78.9 74.9 73.9 80.9 49.4 71.6
[…]
Converting tokenizers to byte-level. As per Section 3.2, we need a procedure to convert tokenizers to the byte level to ensure that token decomposition is always possible. This is trivial in most cases; the missing bytes just need to be added to the vocabulary. BPE is an exception: here, we need to change the units on which merges are defined from characters to bytes. We achieve this by adding merges to assemble the characters used by the tokenizer from their constituent bytes to the beginning of the merge table. This preserves the tokenization in more than 99% of cases (Appendix  J).
Converting tokenizers to UnigramLM. We also introduce a procedure to convert arbitrary tokenizers to tokenizers using UnigramLM as the tokenization function. We refer to this process as unigramifying (details in Appendix A). An important assumption of the hypernetwork training is that by using the UnigramLM parametrization with scores distributed as Gaussians we can cover a sufficiently diverse distribution of tokenizers for the hypernetwork to generalize to e.g. BPE tokenizers. Unigramifying allows us to check if, in principle, this is possible. Luckily, we find that it is: unigramifying results in minimal performance degradation when substituting the original tokenizer with the corresponding UnigramLM tokenizer (Appendix J). Although this does not guarantee that our distribution of tokenizers is sufficiently diverse, our empirical results suggest it is (cf. Section 4.2).
@@
What is the effect of amortizing over the tokenization function? As described earlier in Section 3, we ‘amortize’ over the tokenization function, that is, the tokenization function is not an input to our hypernetwork. We find that the predicted amortized embeddings are robust to the choice of tokenization function. For example, the set of embeddings predicted for the GPT2 vocabulary has low bits-per-character for both the original GPT2 tokenization function and a different UnigramLM tokenization function with scores based on token frequencies (Appendix J). This is not the case for the original GPT2 embeddings: while they (as expected) perform well with the original GPT2 tokenizer, there is significant performance degradation when switching to the frequency-based UnigramLM tokenization function. This calls into question prior work copying the embeddings of overlapping tokens for transfer across tokenizers (Dobler & de Melo, 2023; Gee et al., 2022, among others), indicating that even if there is an exactly overlapping token in the original tokenizer, it is not necessarily the optimal initialization of the corresponding token in the new tokenizer.
@@
Although we amortize over most of the aspects of the tokenization function, in practice, tokenization functions rely on a considerable amount of engineering, so it is not possible to amortize over everything; we discuss remaining assumptions in Appendix I.
@@
Analyzing computational overhead. We estimate the FLOPs per token of multiple hypernetworks in Appendix K. Given a batch size $n$ and sequence length $s$ for the main model, and using the hypernetwork to compose $k$ token sequences of length $t$, the FLOPs per batch will be $n\times s\times(\frac{\text{FLOPs}}{\text{token}})_{\text{main}}+k\times t\times(\frac{\text{FLOPs}}{\text{token}})_{\text{hypernet}}$. Taking Mistral-7B as an example with $n=s=128$, $k=32768$ and $t=7$ the FLOPs per batch will be 252T + 30T i.e. a 12% overhead from applying the hypernet. Notably, we observed that a hypernetwork size of three layers is sufficient, regardless of the main model, so the relative overhead decreases with increased amounts of layers in the main model.
Generalization to unseen tokens. Although our primary goal is generalization to unseen tokenizers (i.e., tuples $(\mathcal{V},T)$), the question of how well our hypernetwork can generalize to unseen tokens (elements of $\mathcal{V}$) presents itself. To answer this question, we test the XLM-R and Mistral-7B hypernetworks on out-of-distribution vocabularies. Specifically, we test the XLM-R hypernetwork on Farsi and Dutch (which are unseen by the hypernet, but seen by the base model) as well as Aymara and Guarani, which are unseen by both. Table 6 confirms the hypernet performs well in this case, even gaining in performance over the model with original embeddings in completely unseen languages. In this setup, up to 40% of the used tokens in the target vocabularies have never been seen during hypernetwork training (we analyze this overlap in detail in Appendix G). The reason for the performance increase from the hypernetwork on unseen languages may be that, under the original tokenization, the embeddings of many tokens occuring in unseen languages are undertrained (c.f. Land & Bartolo, 2024), while the embeddings produced by the hypernetwork do not suffer from this issue; future work could investigate this in more detail. For Mistral-7B, we instead transfer to an out-of-distribution word-level tokenizer by creating a tokenizer which contains all words which occur in any evaluation corpus (approx. 100k in total). 3.3k words are completely unseen and 13.5k words have been seen in less than 0.1% of training steps. Still, performance only deteriorates by a small amount and the improvement over FOCUS persists as shown in Table 7.